Question # 1: A retailer has yearly sales of $650,000. Inventory on January 1 is $250,000 (at cost). During the year, $500,000 of merchandise (at cost) is purchased. The ending inventory is $275,000 (at cost). Operating costs are $90,000.

a. Calculate the cost of goods sold

b. Calculate the net profit
Answer #1

1.
Cost of Goods Sold=Beginning Inventory+Purchases-Ending Inventory=250000+500000-275000=475000

2.
=Sales-Cost of Goods Sold-Operating Costs
=650000-475000-90000
=85000